# Smooth Lower Bounds for Differentially Private Algorithms via Padding-and-Permuting Fingerprinting Codes

### v3

@@ Appendix A Fingerprinting Codes
  • [1] Input: Number of users $n\in{\mathbb{N}}$ and a confidence parameter $\beta\in[0,1]$.
  • [2] Operation:  1. Let $d\geq\Theta(n^{2}\log^{2}(n)\log(1/\beta))$ be the length of the code. 2. Sample $(x_{1},\ldots,x_{n},z)\sim{\cal D}^{\prime}(n,d)$ (Section 4.2). 3. Output $X=(x_{i}^{j})_{i\in[n],j\in[d]}$ and share $z=(z^{1},\ldots,z^{d})$ with $\mathsf{Trace}$.
@@
  • [3] Input: A codebook $X\in\{-1,1\}^{n\times d}$, a word $q\in\{-1,1\}^{d}$, and a shared state $z\in\{-1,1\}^{d}$.
  • [4] Operation:  1. For every $i\in[n]$: If $\langle x_{i},q\rangle-\langle z,q\rangle>\frac{0.2d}{n\cdot\ln(5n)}$, output $i$. 2. Else (no such $i\in[n]$), output $\perp$.

### v4

@@ Appendix A Fingerprinting Codes
Input: Number of users $n\in{\mathbb{N}}$ and a confidence parameter $\beta\in[0,1]$. Operation: 1. Let $d\geq\Theta(n^{2}\log^{2}(n)\log(1/\beta))$ be the length of the code. 2. Sample $(x_{1},\ldots,x_{n},z)\sim{\cal D}^{\prime}(n,d)$ (Section 4.2). 3. Output $X=(x_{i}^{j})_{i\in[n],j\in[d]}$ and share $z=(z^{1},\ldots,z^{d})$ with $\mathsf{Trace}$.
@@
Input: A codebook $X\in\{-1,1\}^{n\times d}$, a word $q\in\{-1,1\}^{d}$, and a shared state $z\in\{-1,1\}^{d}$. Operation: 1. For every $i\in[n]$: If $\langle x_{i},q\rangle-\langle z,q\rangle>\frac{0.2d}{n\cdot\ln(5n)}$, output $i$. 2. Else (no such $i\in[n]$), output $\perp$.